# An Actionable Framework for Assessing Bias and Fairness in Large Language Model Use Cases

### v2

@@ 2.3 Mapping Bias and Fairness Risks to LLM Use Cases
First, consider use cases where an LLM generates text outputs that are not constrained to a predefined set of classes (e.g., positive vs. negative) or list elements (e.g., products to recommend). For the sake of brevity, this group of use cases will be hereafter referred to as “text generation and summarization," acknowledging that this category can encompass additional use cases including, but not limited to, machine translation, retrieval augmented generation (RAG), and question-answering. An example of a text generation use case could be utilizing an LLM to compose personalized messages for customer outreach. Similarly, a summarization use case may involve employing an LLM to extract pertinent information or provide summaries from clinical notes. These use cases carry the potential risk of generating toxic text in their outputs. Moreover, if these use cases fail to meet the criteria of FTU, meaning that the prompts include references to protected attributes, they also pose the risk of perpetuating stereotypes or exhibiting counterfactual unfairness.
Classification.
@@
LLMs have been widely used for text classification (Sun et al., 2023; Widmann and Wich, 2022; Bonikowski et al., 2022; Howard and Ruder, 2018; Sun et al., 2019; Chai et al., 2020; Chen et al., 2020; Lin et al., 2021). In the context of bias and fairness, it is important to distinguish whether the text inputs can be mapped to a protected attribute, either by containing direct mentions of a protected attribute group, or in the case of person-level prompt granularity, corresponding to individuals belonging to certain protected attribute groups. For instance, one example of a person-level classification use case could involve utilizing an LLM to classify customer feedback as positive or negative in order to assign appropriate follow-ups. Similar to traditional person-level classification challenges in machine learning, these use cases present the risk of allocational harms. On the other hand, classification use cases that do not involve person-level data and satisfy FTU are not subject to these bias and fairness risks.
Recommendation.
@@
Recommendation is another potential application of LLMs (Bao et al., 2023; Gao et al., 2023), such as using an LLM to recommend products to customers. Zhang et al. (2023) show that LLMs used as recommendation engines can discriminate when exposed to protected attribute information. Given this concern, it follows that LLM recommendation use cases pose the risk of counterfactual unfairness if they do not satisfy FTU.
@@ 3 Bias and Fairness Evaluation Metrics
This section introduces a range of evaluation metrics segmented by the use case task and its applicable bias and fairness risks. The proposed framework encompasses three distinct use case categories: 1) text generation and summarization, 2) classification, and 3) recommendation. For each category, bias and fairness evaluation metrics are included that address the applicable risks.
Accounting for prompt-specific risk is important to accurately reflect the risk of a specific use case, as Wang et al. (2024b) find in their evaluations that toxicity probability is 26 to 101 times higher when dealing with toxic prompts compared to non-toxic prompts. Moreover, Goldfarb-Tarrant et al. (2020); Delobelle et al. (2021) have demonstrated that evaluation metrics which take into account the LLM’s task provide a more accurate reflection of the associated risk compared to metrics based on embeddings or token probabilities. Accordingly, the objective of this work is to assess the risks associated with a particular use case, taking into consideration not only the LLM used but also the task at hand and the prompt population. Following the characterization of a use case based on a model and a known population of prompts, as established in Section 2.1, each metric definition presented below is contextualized within an evaluation sample of size $N$ drawn from a known population of prompts $\mathcal{P}_{X}$.
3.1 Bias Metrics for Text Generation and Summarization Use Cases
Gallegos et al. (2023) propose a detailed taxonomy of evaluation metrics for bias evaluations in LLMs. They partition these metrics into three main categories: embedding-based metrics, probability-based metrics, and generated-text metrics. While embedding- and probability-based metrics require access to an LLM’s upstream architecture, generated-text metrics instead treat an LLM like a black box and can be easily calculated from LLM output alone. Hence, given their ease of application, only generated-text metrics are included in the framework proposed here.
@@ 3.1 Bias Metrics for Text Generation and Summarization Use Cases
While Gallegos et al. (2023) further segment generated-text metrics into three subcategories (namely distribution metrics, classifier metrics, and lexicon metrics), this work segments these metrics according to the risk taxonomy outlined in Section 2.2. In particular, generated-text bias and fairness metrics are segmented into the following categories: toxicity metrics, stereotype metrics, and counterfactual fairness metrics. Toxicity metrics leverage a pre-trained toxicity classifier, such as Perspective API [^12] https://perspectiveapi.com, to assign a toxicity score to an LLM’s output (Chowdhery et al., 2022; Lees et al., 2022; Wang et al., 2024b; Bommasani et al., 2023; Gehman et al., 2020). Stereotype metrics assess the relative co-occurrence of stereotype words with protected attribute words (Bordia and Bowman, 2019; Bommasani et al., 2023) or leverage a pre-trained stereotype classifier to assign a stereotype score to an LLM’s output (Zekun et al., 2023). Lastly, counterfactual fairness metrics assess fairness through a causal lens by exploiting variation in mentions of protected attribute groups in otherwise-identical prompts (Huang et al., 2020). Select metrics are detailed in each category below. [^13] Note that the discussion of generated text bias metrics in this paper is not exhaustive. Rather, the intention is to select an appropriate set of evaluation metrics to address each risk to be assessed. Notable metrics omitted from this framework include HONEST (Nozza et al., 2021), score parity (Sicilia and Alikhani, 2023), Full Gen Bias (Smith et al., 2022), regard score (Sheng et al., 2019), demographic representations (Bommasani et al., 2023), psycholinguistic norms (Dhamala et al., 2021), and gender polarity (Dhamala et al., 2021).
@@ 3.1.1 Toxicity Metrics
Toxicity metrics leverage a pre-trained toxicity classifier to assess general toxicity that is not specific to a particular protected attribute. Given a toxicity classifier $T:\mathcal{Y}\xrightarrow[]{}[0,1]$, this section defines toxicity metrics in the context of an LLM use case $(\mathcal{M},\mathcal{P}_{X})$. Specifically, these metrics are evaluated on a set of $m$ responses per prompt, $\{\hat{Y}_{ij}\}_{j=1}^{m}$, generated by $\mathcal{M}$, from a sample of prompts $X_{1},...,X_{N}$ drawn from $\mathcal{P}_{X}$. In practice, variation in responses for the same prompt can be achieved by choosing a non-zero temperature for the LLM or by utilizing various random seeds (Wang et al., 2024b). Below, formal definitions of three toxicity metrics are provided: Expected Maximum Toxicity, Toxicity Probability, and Toxic Fraction. All three metrics range in values from 0 to 1, with smaller values indicating a greater degree of fairness.
@@
Toxicity Fraction (TF) (Bommasani et al., 2023).
@@
This metric effectively estimates the likelihood that responses generated by $\mathcal{M}$ on prompts from $\mathcal{P}_{X}$ contain toxic text (Bommasani et al., 2023). Note that while the standard choice of $m$ for this metric is $m=1$ (Bommasani et al., 2023), a larger value of $m$ may be preferred in practice if sampling a large $N$ is infeasible.
@@ 3.1.2 Stereotype Metrics
Stereotype metrics aim to identify harmful stereotypes specific to protected attributes that might be present in an LLM’s output. Because these metrics rely on mentions of protected attribute groups, these metrics may be unnecessary if FTU is satisfied for an LLM use case. [^14] Note that stereotype risk, while low, may still exist even if FTU is satisfied for an LLM use case. Among stereotype metrics, this work distinguishes between metrics based on co-occurrence of protected attribute words and stereotypical words, and metrics that leverage a stereotype classifier.
@@ 3.1.2.1 Co-occurrence-Based Metrics
This section outlines a set of metrics that assess stereotype risk based on relative co-occurrence of protected attribute words with neutral words of interest. In particular, formal definitions of two co-occurrence-based stereotype metrics are provided: Co-Occurrence Bias Score and Stereotypical Associations.
@@
Stereotypical Associations (SA) (Bommasani et al., 2023).
Consider a set of protected attribute groups $\mathcal{G}$, an associated set of protected attribute group lexicons $\mathcal{A}$, and an associated set of stereotypical words $W$. Additionally, let $C(x,\hat{Y})$ denote the number of times that the word $x$ appears in the output $\hat{Y}$, $P^{\text{ref}}$ denote a reference distribution, and $TVD$ denote total variation difference. [^16] The reference distribution recommended by Bommasani et al. (2023) is the uniform distribution. Total variation distance measures the distance between probability distributions. For a given LLM $\mathcal{M}(X;\theta)$ and a sample of prompts $X_{1},...,X_{N}$ drawn from $\mathcal{P}_{X}$, the full computation of SA is as follows:
@@
In words, SA measures the relative co-occurrence of a set of stereotypically associated words across protected attribute groups. SA ranges in value from 0 to 1, where smaller values indicate greater fairness.
@@ 3.1.2.2 Metrics Leveraging a Stereotype Classifier
It has been shown that stereotype classifiers can be an effective tool for assessing stereotype risk in LLM use cases (Zekun et al., 2023). [^17] Zekun et al. (2023) refer to https://huggingface.co/Narrativa/distilroberta-finetuned-stereotype-detection. This section introduces three new metrics by extending the toxicity metrics outlined in Section 3.1.1, leveraging a pre-trained stereotype classifier $St:\mathcal{Y}\xrightarrow[]{}[0,1]$ rather than a toxicity classifier. Namely, these metrics include: Expected Maximum Stereotype, Stereotype Probability, and Stereotype Fraction. All three metrics range in values from 0 to 1, with smaller values indicating a greater degree of fairness.
@@
effectively estimating the likelihood that responses generated by $\mathcal{M}$ on prompts from $\mathcal{P}_{X}$ contain stereotypes. Note that while the standard choice of $m$ for the analogous toxicity metric, EMT, is $m=1$ (Bommasani et al., 2023), a larger value of $m$ may be preferred in practice if sampling a large $N$ is infeasible.
@@ 3.1.3 Counterfactual Fairness Metrics
Counterfactual metrics aim to assess differences in LLM output when different protected attributes are mentioned in input prompts, all else equal. Given two protected attribute groups $G^{\prime},G^{\prime\prime}$, these metrics are defined in the context of an LLM use case $(\mathcal{M},\mathcal{P}_{X})$. In particular, these metrics are evaluated on a sample of counterfactual response pairs $(\hat{Y}_{1}^{\prime},\hat{Y}_{1}^{\prime\prime}),...,(\hat{Y}_{N}^{\prime},\hat{Y}_{N}^{\prime\prime})$ generated by $\mathcal{M}$, from a sample of counterfactual input pairs $(X_{1}^{\prime},X_{1}^{\prime\prime}),...,(X_{N}^{\prime},X_{N}^{\prime\prime})$ drawn from $\mathcal{P}_{X|\mathcal{A}}$. [^18] In practice, counterfactual substitution can be achieved by leveraging a mapping of one protected attribute group lexicon to another. For instance, a female to male lexicon mapping could include substitutions such as ’she’: ’he’, ’hers’: ’his’, ’her’: ’him’, ’herself’: ’himself’, ’female’: ’male’, ’females’: ’males’, ’woman’: ’man’, ’women’: ’men’, ’girl’: ’boy’, ’girls’: ’boys’, ’daughter’: ’son’, ’daughters’: ’sons’, ’mother’: ’father’, ’mothers’: ’fathers’, ’sister’: ’brother’, ’sisters’: ’brothers’, ’aunt’: ’uncle’, ’aunts’: ’uncles’, ’niece’: ’nephew’, ’nieces’: ’nephews’, ’lady’: ’gentleman’, ’ladies’: ’gentlemen’, ’grandmother’: ’grandfather’, ’grandmothers’: ’grandfathers’. It is important to note that the mapping does not need to be exhaustive, as long as it has sufficient coverage to generate a large sample of counterfactual input pairs. Note that, in scenarios where a large $N$ is infeasible, practitioners may opt to generate multiple response pairs per counterfactual input pair.
These metrics, which are categorized into counterfactual similarity metrics and counterfactual sentiment metrics, respectively quantify the differences in text similarity and sentiment by leveraging the variations in LLM output observed across counterfactual input pairs. Due to their reliance on mentions of protected attributes in input prompts, if FTU is satisfied for an LLM use case, these metrics need not be used.
@@ 3.1.3.1 Counterfactual Similarity
Counterfactual similarity metrics measure the similarity in outputs generated from counterfactual input pairs according to a specified invariance metric $\upsilon$, i.e. $\upsilon(\mathcal{M}(X^{\prime};\theta),\mathcal{M}(X^{\prime\prime};\theta)).$ These metrics effectively assess whether the LLM use case satisfies the counterfactual invariance property defined in Section 2.2. One such example of $\upsilon$ is exact match (Rajpurkar et al., 2016), but (Gallegos et al., 2023) argue that this metric is too strict. This work introduces three, less stringent, counterfactual similarity metrics: Counterfactual ROUGE-L, Counterfactual BLEU, and Counterfactual Cosine Similarity, which are extensions of state-of-the-art text similarity metrics (Minaee et al., 2024; Lin, 2004; Papineni et al., 2002; Singhal and Google, 2001; Gomaa and Fahmy, 2013). While the first two assess similarity using token-sequence overlap, the the third assesses similarity using sentence embeddings. All three metrics range in values from 0 to 1, with larger values indicating a greater degree of fairness.
@@
This work introduces CROUGE-L, defined as the average ROUGE-L score (Lin, 2004) over counterfactually generated output pairs. The full calculation of CROUGE-L is as follows:
@@
This work introduces CBLEU, defined as the average BLEU score (Papineni et al., 2002) over counterfactually generated output pairs. The full calculation of CBLEU is as follows:
@@
where $snt$ denotes a sentence in an LLM output, $len(\hat{Y})$ denotes the number of tokens in an LLM output, $C(b{\text{-}}gram,\hat{Y}_{i}^{\prime})$ denotes the number of times $b{\text{-}}gram$ appears in $\hat{Y}_{i}^{\prime}$ and $C(b{\text{-}}gram,\hat{Y}_{i}^{\prime}|\hat{Y}_{i}^{\prime\prime})$ denotes the number of times $b{\text{-}}gram$ appears in $\hat{Y}_{i}^{\prime}$ given that it also appears in $\hat{Y}_{i}^{\prime\prime}$ (Papineni et al., 2002; goo, ). For symmetry, the minimum of these two BLEU scores for each counterfactual pair is obtained before averaging. For the same reasons as with CROUGE-L, practitioners should mask protected attribute words in counterfactual output pairs before computing CBLEU.
@@
Given a sentence transformer $\mathbf{V}:\mathcal{Y}\xrightarrow{}\mathbb{R}^{d}$, this work defines CCS as:

### v3

@@ 2.3.1 Text Generation and Summarization
We first consider use cases where an LLM generates text outputs that are not constrained to a predefined set of classes (e.g., positive vs. negative) or list elements (e.g., products to recommend). For the sake of brevity, we will hereafter refer to this group of use cases as “text generation and summarization," acknowledging that this category can encompass additional tasks including machine translation, question-answering, and others. One example of this type of use case is using an LLM to compose personalized messages for customer outreach. Use cases in this category carry the risk of generating toxic text in their outputs. Moreover, if these use cases fail to satisfy FTU, meaning that the prompts include mentions of protected attributes, they also pose the risk of perpetuating stereotypes or exhibiting counterfactual unfairness.
2.3.2 Classification
@@
LLMs have been widely used for text classification (Sun et al., 2023; Widmann and Wich, 2022; Bonikowski et al., 2022; Howard and Ruder, 2018; Sun et al., 2019; Chai et al., 2020; Chen et al., 2020; Lin et al., 2021). In the context of bias and fairness, it is important to distinguish whether the text inputs can be mapped to a protected attribute, either by containing direct mentions of a protected attribute group, or in the case of person-level prompt granularity, corresponding to individuals belonging to certain protected attribute groups. For instance, using an LLM to classify customer feedback as positive or negative in order to assign appropriate follow-ups would be an example of a person-level classification use case. Similar to traditional person-level classification problems in machine learning, these use cases present the risk of allocational harms. On the other hand, classification use cases that do not involve person-level data and satisfy FTU are not subject to these bias and fairness risks.
2.3.3 Recommendation
@@
Recommendation is another potential application of LLMs (Bao et al., 2023; Gao et al., 2023), such as using an LLM to recommend products to customers. Zhang et al. (2023) show that LLMs used as recommendation engines can discriminate when exposed to protected attribute information. It follows that LLM recommendation use cases pose the risk of counterfactual unfairness if they do not satisfy FTU.
@@ 3 Bias and Fairness Evaluation Metrics
Our proposed framework encompasses three distinct use case categories: 1) text generation and summarization, 2) classification, and 3) recommendation. For each category, we present various evaluation metrics that address the applicable bias and fairness risks. For practical reasons, we limit the selection of metrics to those requiring only LLM generated output for computation. [^14] Due to practical limitations of the input requirements, we omit from the framework both embedding-based metrics, which are computed using an LLM’s hidden vector representations of words or sentences (Islam et al., 2016; May et al., 2019; Guo and Caliskan, 2020), and probability-based metrics, which leverage predicted token probabilities from an LLM (Webster et al., 2020; Kurita et al., 2019; Ahn and Oh, 2021; Kaneko and Bollegala, 2021; Salazar et al., 2019; Barikeri et al., 2021; Nangia et al., 2020; Nadeem et al., 2021). While these metrics require access to an LLM’s upstream architecture, the metrics we include in our framework do not. Importantly, we note that metrics focused on the downstream task, consistent with the metrics incorporated in this framework, have been shown to be more reliable than metrics derived from embeddings or token probabilities (Goldfarb-Tarrant et al., 2020; Delobelle et al., 2021). To ensure our metric definitions accurately reflect use-case-specific nature of our framework, we contextualize each metric within an evaluation sample of size $N$ drawn from a known population of prompts $\mathcal{P}_{X}$.
3.1 Metrics for Text Generation and Summarization Use Cases
@@
We segment metrics for text generation and summarization use cases based on the applicable bias and fairness risks, as outlined in Section 2.2. Namely, this includes toxicity metrics, stereotype metrics, and counterfactual fairness metrics. Toxicity metrics leverage a pre-trained toxicity classifier, such as Perspective API [^15] https://perspectiveapi.com, to assign a toxicity score to an LLM’s output (Chowdhery et al., 2022; Lees et al., 2022; Wang et al., 2024b; Liang et al., 2023; Gehman et al., 2020). Stereotype metrics assess the relative co-occurrence of stereotype words with protected attribute words (Bordia and Bowman, 2019; Liang et al., 2023) or leverage a pre-trained stereotype classifier to assign a stereotype score to an LLM’s output (Zekun et al., 2023). Lastly, counterfactual fairness metrics assess fairness through a causal lens by exploiting variation in mentions of protected attribute groups in otherwise-identical prompts (Huang et al., 2020). Select metrics are detailed in each category below. [^16] We acknowledge that our set of bias and fairness metrics is not exhaustive. Rather, the intention is to select an appropriate set of evaluation metrics to address each risk to be assessed. Notable metrics omitted from this framework include HONEST (Nozza et al., 2021), score parity (Sicilia and Alikhani, 2023), Full Gen Bias (Smith et al., 2022), regard score (Sheng et al., 2019), demographic representations (Liang et al., 2023), psycholinguistic norms (Dhamala et al., 2021), and gender polarity (Dhamala et al., 2021).
@@ 3.1.1 Toxicity Metrics
Toxicity metrics leverage a pre-trained toxicity classifier, $T:\mathcal{Y}\xrightarrow[]{}[0,1]$, that maps a text input to a toxicity score. Here, we define toxicity metrics in the context of an LLM use case $(\mathcal{M},\mathcal{P}_{X})$. Specifically, these metrics are computed on a set of $m$ responses per prompt, $\{\hat{Y}_{ij}\}_{j=1}^{m}$, generated by $\mathcal{M}$, from a sample of prompts $X_{1},...,X_{N}$ drawn from $\mathcal{P}_{X}$. In practice, variation in responses for the same prompt can be achieved by choosing a non-zero temperature for the LLM or by utilizing various random seeds (Wang et al., 2024b). Below, we define three toxicity metrics: Expected Maximum Toxicity, Toxicity Probability, and Toxic Fraction. All three metrics range in value from 0 to 1, with smaller values indicating a greater degree of fairness.
@@
Toxicity Fraction (TF) (Liang et al., 2023).
@@
This metric effectively estimates the likelihood that responses generated by $\mathcal{M}$ on prompts from $\mathcal{P}_{X}$ contain toxic text (Liang et al., 2023). Note that while the standard choice of $m$ for this metric is $m=1$ (Liang et al., 2023), a larger value of $m$ may be preferred in practice if sampling a large $N$ is infeasible.
@@ 3.1.2 Stereotype Metrics
Stereotype metrics aim to identify harmful stereotypes specific to protected attributes that might be present in an LLM’s output. Because these metrics rely on mentions of protected attribute groups, these metrics may be unnecessary if FTU is satisfied for an LLM use case. [^17] Note that stereotype risk, while low, may still exist even if FTU is satisfied for an LLM use case. Among stereotype metrics, we distinguish between metrics based on co-occurrence of protected attribute words and stereotypical words, and metrics that leverage a stereotype classifier.
@@ 3.1.2.1 Co-occurrence-Based Metrics
In this section, we outline a set of metrics that assess stereotype risk based on relative co-occurrence of protected attribute words with stereotype words of interest. These metrics effectively assess the degree to which equal group associations, as defined in Section 2, is satisfied. We define two co-occurrence-based stereotype metrics: Co-Occurrence Bias Score and Stereotypical Associations.
@@
Stereotypical Associations (SA) (Liang et al., 2023).
Consider a set of protected attribute groups $\mathcal{G}$, an associated set of protected attribute group lexicons $\mathcal{A}$, and an associated set of stereotypical words $W$. Additionally, let $C(x,\hat{Y})$ denote the number of times that the word $x$ appears in the output $\hat{Y}$, $P^{\text{ref}}$ denote a reference distribution, and $TVD$ denote total variation difference. [^19] The reference distribution recommended by Liang et al. (2023) is the uniform distribution. Total variation distance measures the distance between probability distributions. For a given LLM $\mathcal{M}(X;\theta)$ and a sample of prompts $X_{1},...,X_{N}$ drawn from $\mathcal{P}_{X}$, the full computation of SA is as follows:
@@
In words, SA measures the relative co-occurrence of a set of stereotypically associated words across protected attribute groups. [^20] Note that while COBS and SA both assess equal group associations, COBS is computed pairwise, while SA is computed attribute-wise. SA ranges in value from 0 to 1, where smaller values indicate greater fairness.
@@ 3.1.2.2 Metrics Leveraging a Stereotype Classifier
It has been shown that stereotype classifiers can be an effective tool for assessing stereotype risk in LLM use cases (Zekun et al., 2023). [^21] Zekun et al. (2023) use https://huggingface.co/Narrativa/distilroberta-finetuned-stereotype-detection for their classifier. We introduce three stereotype metrics by extending the toxicity metrics outlined in Section 3.1.1, leveraging a pre-trained stereotype classifier, $St:\mathcal{Y}\xrightarrow[]{}[0,1]$, rather than a toxicity classifier. Namely, these metrics include: Expected Maximum Stereotype, Stereotype Probability, and Stereotype Fraction. All three metrics range in value from 0 to 1, with smaller values indicating a greater degree of fairness. We define these metrics below.
@@
effectively estimating the likelihood that responses generated by $\mathcal{M}$ on prompts from $\mathcal{P}_{X}$ contain stereotypes. Note that while the standard choice of $m$ for the analogous toxicity metric, EMT, is $m=1$ (Liang et al., 2023), a larger value of $m$ may be preferred in practice if sampling a large $N$ is infeasible.
@@ 3.1.3 Counterfactual Fairness Metrics
Counterfactual metrics aim to assess differences in LLM output when different protected attributes are mentioned in input prompts, all else equal. Given two protected attribute groups $G^{\prime},G^{\prime\prime}$, we define these metrics in the context of an LLM use case $(\mathcal{M},\mathcal{P}_{X})$. In particular, these metrics are evaluated on a sample of counterfactual response pairs $(\hat{Y}_{1}^{\prime},\hat{Y}_{1}^{\prime\prime}),...,(\hat{Y}_{N}^{\prime},\hat{Y}_{N}^{\prime\prime})$ generated by $\mathcal{M}$, from a sample of counterfactual input pairs $(X_{1}^{\prime},X_{1}^{\prime\prime}),...,(X_{N}^{\prime},X_{N}^{\prime\prime})$ drawn from $\mathcal{P}_{X|\mathcal{A}}$. [^22] In practice, counterfactual substitution can be achieved by leveraging a mapping of one protected attribute group lexicon to another. For instance, a female to male lexicon mapping could include substitutions such as ’she’: ’he’, ’hers’: ’his’, ’her’: ’him’, ’herself’: ’himself’, ’female’: ’male’, ’females’: ’males’, ’woman’: ’man’, ’women’: ’men’, ’girl’: ’boy’, ’girls’: ’boys’, ’daughter’: ’son’, ’daughters’: ’sons’, ’mother’: ’father’, ’mothers’: ’fathers’, ’sister’: ’brother’, ’sisters’: ’brothers’, ’aunt’: ’uncle’, ’aunts’: ’uncles’, ’niece’: ’nephew’, ’nieces’: ’nephews’, ’lady’: ’gentleman’, ’ladies’: ’gentlemen’, ’grandmother’: ’grandfather’, ’grandmothers’: ’grandfathers’. It is important to note that the mapping does not need to be exhaustive, as long as it has sufficient coverage to generate a large sample of counterfactual input pairs. Note that, in scenarios where a large $N$ is infeasible, practitioners may opt to generate multiple response pairs per counterfactual input pair, as is done for toxicity metrics in Section 3.1.1.
These metrics, which we categorize into counterfactual similarity metrics and counterfactual sentiment metrics, respectively quantify the differences in text similarity and sentiment by leveraging the variations in LLM output observed across counterfactual input pairs. Due to their reliance on mentions of protected attributes in input prompts, if FTU is satisfied for an LLM use case, these metrics need not be used.
@@ 3.1.3.1 Counterfactual Similarity
Counterfactual similarity metrics measure the similarity in outputs generated from counterfactual input pairs according to a specified invariance metric $\upsilon$, i.e. $\upsilon(\mathcal{M}(X^{\prime};\theta),\mathcal{M}(X^{\prime\prime};\theta)).$ These metrics effectively assess whether the LLM use case satisfies the counterfactual invariance property defined in Section 2.2. One such example of $\upsilon$ is exact match (Rajpurkar et al., 2016), but Gallegos et al. (2023) argue that this metric is too strict. We introduce three, less stringent, counterfactual similarity metrics: Counterfactual ROUGE-L, Counterfactual BLEU, and Counterfactual Cosine Similarity, which are extensions of state-of-the-art text similarity metrics (Minaee et al., 2024; Lin, 2004; Papineni et al., 2002; Singhal and Google, 2001; Gomaa and Fahmy, 2013). The first two assess similarity using token-sequence overlap and range in value from 0 to 1. The third assesses similarity using sentence embeddings and ranges in value from -1 to 1. For each, larger values indicate a greater degree of fairness.
@@
We introduce CROUGE-L, defined as the average ROUGE-L score (Lin, 2004) over counterfactually generated output pairs. The full calculation of CROUGE-L is as follows:
@@
We define CBLEU as the average BLEU score (Papineni et al., 2002) over counterfactually generated output pairs. The full calculation of CBLEU is as follows:
@@
where $snt$ denotes a sentence in an LLM output, $len(\hat{Y})$ denotes the number of tokens in an LLM output, $C(b{\text{-}}gram,\hat{Y}_{i}^{\prime})$ denotes the number of times $b{\text{-}}gram$ appears in $\hat{Y}_{i}^{\prime}$ and $C(b{\text{-}}gram,\hat{Y}_{i}^{\prime}|\hat{Y}_{i}^{\prime\prime})$ denotes the number of times $b{\text{-}}gram$ appears in $\hat{Y}_{i}^{\prime}$ given that it also appears in $\hat{Y}_{i}^{\prime\prime}$ (Papineni et al., 2002; goo, ). To achieve symmetry, we use the minimum of these two BLEU scores for each counterfactual pair is obtained before averaging. For the same reasons as with CROUGE-L, practitioners should mask protected attribute words in counterfactual output pairs before computing CBLEU.
@@
Given a sentence transformer $\mathbf{V}:\mathcal{Y}\xrightarrow{}\mathbb{R}^{d}$, we define CCS as: